In a shipment of 22 smart phones, 3 are defective. How many ways can a quality control inspector randomly test 5 smart phones, of which 3 are defective?
I tried doing C(5,3) and P(5,3). Neither are correct. It says the answer is 171. I really don't know what I'm doing wrong. Thanks in advance.

Answer by stanbon(75887)

In a shipment of 22 smart phones, 3 are defective. How many ways can a quality control inspector randomly test 5 smart phones, of which 3 are defective?
I tried doing C(5,3) and P(5,3). Neither are correct. It says the answer is 171. I really don't know what I'm doing wrong. Thanks in advance.
----
If 3 of the 5 are defective, only 2 are selected from the 19 that
are not defective.
Ans: 19C2 = (19*18)/(1*2) = 19*9 = 171
-------------
Cheers,
Stan H.
------------